ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 235, 2020

Adams, T., V. Bergelson and W. Sun Under- and overindependence in measure preserving systems

349

Alon, N., I. Balla, L. Gishboliner, A. Mond and F. Mousset The minrank of random graphs over arbitrary fields

63

An, Y. and H. Liu The Schrödinger–Poisson type system involving a critical nonlinearity on the first Heisenberg group

385

Balla, I. See N. Alon Barry, D., A. Chapman and A. Laghribi The descent of biquater-

nion algebras in characteristic two

295

Bergelson, V. See T. Adams Bernshteyn, A. Ergodic theorems for the shift action and pointwise

versions of the Abért–Weiss Theorem

255

Bloom, T. F. and A. Walker GCD sums and sum-product estimates

1

Buhovsky, L., A. Kiro and S. Sodin Power substitution in quasianalytic Carleman classes

79

Chai, J. and Z. Qi Bessel identities in the Waldspurger correspondence over the complex numbers

439

Chapman, A. See D. Barry Chen, W., S. Garti and T. Weinert Cardinal characteristics of

the continuum and partitions

13

Coulangeon, R. and G. Nebe Slopes of Euclidean lattices, tensor product and group actions

39

De Thélin, H., T. Gauthier and G. Vigny The bifurcation measure has maximal entropy

213

Elekes, M., D. Nagy, M. Poór and Z. Vidnyánszky A Haar meager set that is not strongly Haar meager

91

Garti, S. See W. Chen Gauthier, T. See H. De Thélin Gishboliner, L. See N. Alon

König, J. On number fields with power-free discriminant  413

Kegel, M., J. Schneider and K. Zehmisch Symplectic dynamics and the 3-sphere

245

Kiro, A. See L. Buhovsky Laghribi, A. See D. Barry Li, H. and Z. Rong Null actions and RIM non-open extensions of

strongly proximal actions

139

Liu, H. See Y. An Lodha, Y. Coherent actions by homeomorphisms on the real line or

an interval

183

Lucchini, A. and G. Tracey Finite groups with large Chebotarev invariant

169

Mond, A. See N. Alon

Mossel, E. Gaussian bounds for noise correlation of resilient functions  111

Mousset, F. See N. Alon Nagy, D. See M. Elekes Nebe, G. See R. Coulangeon Poór, M. See M. Elekes Qi, Z. See J. Chai Rong, Z. See H. Li Schneider, J. See M. Kegel

Shumyatsky, P. Profinite groups with pronilpotent centralizers  325

Sodin, S. See L. Buhovsky Sun, W. See T. Adams Tracey, G. See A. Lucchini

Vasey, S. Tameness from two successive good frames  465

Vidnyánszky, Z. See M. Elekes Vigny, G. See H. De Thélin Walker, A. See T. F. Bloom Weinert, T. See W. Chen Zehmisch, K. See M. Kegel